# WhyFlow: Interrogative Debugger for Sensemaking Taint Analysis

### v2

@@ 6. Discussion
Listing 2: A divergent path query flagging branch points where flow diverges to multiple targets. Using the raw three‐arity edge(id,src,dst) input, we declare branch(n) for each $n$ that appears as src in more than one edge(_,n,_) fact.
@@
Improved Accuracy with Visual Queries.
Participants using WhyFlow achieved better accuracy, especially when analyzing multiple flows. In particular, the AffectedSinks and GlobalImpact queries require the analysis of multiple or interconnected flows. Surprisingly, participants were able to answer questions related to “why-not” using both tools. In contrast, participants using the CodeQL visualizer responses frequently had empty submissions due to lack of time, suggesting viewing results one by one a list view is time-consuming. This shows that an interactive, inquiry-based debugger of taint analysis warnings has the potential to improve sensemaking for large result sets, which is known to hinder adoption (johnson2013don, ).
@@
While participants found WhyFlow “less cumbersome” (P2) for analysis involving multiple flows, they offered suggestions for improving WhyFlow. They recommended adding an on-screen legend to clarify colors and dotted edges, and using distinct shapes for different node types (e.g., sources versus sinks) to avoid visual confusion. Additional interactive features, such as “click-to-edit” functionality for enabling or disabling sanitizers were also proposed. This suggests that participants would appreciate real-time feedback through interaction on the graph.
@@
While WhyFlow supports six template questions from Table 1, adding more template queries is easy since WhyFlow’s template questions are Datalog-style queries. As a case study, adding a new template query called ’divergent path’ query to WhyFlow (Listing 2) took under ten minutes. This query identifies the branch points, which are locations where the flow branches into multiple targets. Note that WhyFlow’s queries supports any taint analysis results from CodeQL. More extension queries are available in our replication package (https://github.com/whyflowtaint/WhyFlow/tree/main/data/extension_queries).
@@
WhyFlow supports interrogative and speculative debugging built on a rich history of using logic programming for software comprehension (holt_1998, ; mens_2002, ; hajiyev_2005, ; eichberg_2008, ; guéhéneuc_2008, ; tourwe_2003, ). Soufflé is the state of the art logic query engine. WhyFlow does not compete with underlying taint analysis engines but instead emphasizes sensemaking support. We pay a one-time cost of fact extraction from CodeQL’s result to support WhyFlow’s interactive features.
Threats to Validity.
A threat to validity is the size of our user study. While our user study involved only 12 participants, it employs a within-subject crossover design, thereby reducing variance and requiring fewer participants (vegas2015crossover, ; Gergle_Tan_2014, ). While our study evaluated WhyFlow only against a single baseline (CodeQL) and and a single subject program (Apache Dubbo), our approach is not tied to CodeQL, and can generalize to any taint analysis.
@@ 7. Related Work
Speculative What-If Analysis. Speculative execution allows the investigation of future or alternative actions developers may perform. Brun et al. explored its application for providing fix suggestions (brun2010speculative, ), as well as identifying conflicts during version merging (brun2011proactive, ). Our work is the first to apply it to taint analysis for reasoning about permissible dataflows of sensitive information under different configuration of different dataflow connectivity.
Templated Questions for asking questions. Developers often ask questions during variety of developer tasks (begel2014analyze, ). In particular, developers need support for understanding code reachability (ko2004designing, ), exploring how different vulnerabilities relate and finding similar ones (smith2015questions, ), tracking intermediate states of analysis (do2018debugging, ), reasoning about system-wide implications when making changes (smith2015questions, ). WhyFlow is the first work providing templated questions for investigating and debugging the results of taint analysis.
Modelling 3rd party libraries. Existing work (li2024llm, ; piskachev2019swan_assist, ; Chiang_Li_Zhou_Banerjee_Schaef_Lyu_Nguyen_Tripp_2024, ) automatically infers models of third-party libraries. However, they do not guide users to understand the impact of these models. They do not support end-user debugging and do not allow end-users to ask questions about the impact of modelling choices. While Paralib (yan2022concept, ) allows comparison between multiple choices of libraries, it does not allow reasoning about how incorrect models cause a mismatch between user expectations and the actual tool results.
End-user debugging. End user debugging (burnett2004end, ; kissinger2006supporting, ; grigoreanu2012end, ) is about the problem of investigating the root cause of tool outcomes, and has been applied to spreadsheet debugging, search engine configuration, etc. Similarly, WhyFlow is an end-user debugger for reasoning how taint flows are affected by the configuration of sources, sinks, and models of third-party libraries.
@@ 8. Conclusion
We presented WhyFlow, a tool for end-user debugging for taint analysis. Through speculative analysis, WhyFlow allows developers to ask “why”, “why-not”, and “what-if” questions about the analysis configuration, and is able to visualize multiple, interconnected flows on a graphical view. WhyFlow enables sensemaking of taint analysis and helps users identify root causes of unexpected flows or missing flows. Our user study confirms that WhyFlow helps users provide 21% more correct answers to questions about the taint analysis, while experiencing a lower cognitive load. These results suggest that WhyFlow’s inquiry-based approach empowers developers to analyze the outputs of taint analysis.
9. Data Availability
The replication package and study’s data are available. at https://github.com/whyflowtaint/WhyFlow.

### v3

@@ 6. Discussion
Listing 2: A divergent path query flagging branch points where flow diverges to multiple targets.
@@
Improved Accuracy with Visual Queries. Using WhyFlow, participants achieved better accuracy, especially when analyzing multiple flows. In particular, the AffectedSinks and GlobalImpact queries require the analysis of multiple flows. Surprisingly, participants were able to answer questions related to “why-not” using both tools. Participants using the CodeQL visualizer frequently had empty submissions due to lack of time. This shows that an interactive, inquiry-based debugger has the potential to improve sensemaking for large result sets, which is known to hinder adoption (johnson2013don, ).
@@
While participants found WhyFlow “less cumbersome” (P2) for analysis involving multiple flows, they offered suggestions for improving WhyFlow. They recommended adding an on-screen legend to clarify colors and dotted edges to avoid visual confusion. Additional interactive features, such as “click-to-edit” functionality for enabling or disabling sanitizers were also proposed. This suggests that participants would appreciate real-time feedback through interaction on the graph.
@@
While WhyFlow supports six template questions from Table 1, adding more template queries is easy since WhyFlow’s template questions are Datalog-style queries. As a case study, adding a new template query called ‘divergent path’ query to WhyFlow (Listing 2) took under ten minutes. This query identifies the branch points, which are locations where the flow branches into multiple targets. Note that WhyFlow’s queries support any taint analysis results from CodeQL.
@@
WhyFlow supports interrogative and speculative debugging built on a rich history of using logic programming for software comprehension (holt_1998, ; mens_2002, ; hajiyev_2005, ; eichberg_2008, ; guéhéneuc_2008, ; tourwe_2003, ). WhyFlow does not compete with underlying taint analysis engines but instead emphasizes sensemaking support.
7. Threats to Validity
Construct Validity.
We designed questions based on realistic “why,” “why-not,” and “what-if” debugging scenarios (ko2004designing, ), requiring reasoning about multiple flows rather than simple lookup tasks. We employed the NASA-TLX questionnaire (hart1988development, ) and the technology acceptance model (lee2003technology, ). However, the self-reported metrics may be influenced by subjective interpretations.
Internal Validity.
We employed a within-subject crossover design with counterbalanced tool order and task assignment to mitigate learning effects and order bias (vegas2015crossover, ). However, participants may experience fatigue, and time pressure. The counterbalanced assignment controls for difficulty differences between problem sets, but residual variations could affect results.
External Validity.
Although our study included only 12 participants, this aligns with typical sample sizes in within-subject SE and HCI studies (vegas2015crossover, ; garcia_2024, ; kang_2024, ; ganji_2023, ; horvath_2022, ; suh_2023, ; huh_2024, ), which require fewer participants than between-subject designs. Most participants were students, but prior work shows that students perform comparably to professionals on security tasks (sandoval_2023, ; acar_2016, ; acar_2017, ; salman_2015, ; naiakshina_2017, ; naiakshina_2020, ; ko_2015, ), and a follow-up study with two security professionals (Section LABEL:casestudy) further corroborated our findings.
While our study evaluated WhyFlow only against CodeQL on Apache Dubbo, WhyFlow operates on generic dataflow facts (nodes, edges, sources, sinks) extractable from any analyzer, template queries are Datalog rules independent of the analysis engine, and the approach is extensible without core modifications (Section 6).
@@ 8. Related Work
Speculative What-If Analysis. Speculative execution (brun2010speculative, ; brun2011proactive, ) allows the investigation of future or alternative actions developers may perform, such as for providing fix suggestions (brun2010speculative, ) and preventing merge conflicts (brun2011proactive, ). Our work is the first to apply it for end-user debugging of taint analysis.
Templated Questions. Developers often ask questions during a variety of developer tasks (begel2014analyze, ). In particular, developers need support for understanding code reachability (ko2004designing, ), exploring how different vulnerabilities relate and finding similar ones (smith2015questions, ), tracking intermediate states of analysis (do2018debugging, ), reasoning about system-wide implications when making changes (smith2015questions, ). WhyFlow is the first work providing templated questions for investigating and debugging the results of taint analysis.
Modelling Third-Party Libraries. Existing work (li2024llm, ; piskachev2019swan_assist, ; Chiang_Li_Zhou_Banerjee_Schaef_Lyu_Nguyen_Tripp_2024, ) automatically infers models of third-party libraries. However, they do not guide users to understand the impact of these models. While Paralib (yan2022concept, ) allows comparison between multiple library choices, it does not allow reasoning about how incorrect models cause a mismatch between user expectations and the actual tool results.
End-user debugging. End user debugging (burnett2004end, ; kissinger2006supporting, ; grigoreanu2012end, ) focuses on investigating the root causes of tool outcomes, and has been applied to spreadsheet debugging, etc. WhyFlow is an end-user debugger for reasoning how taint flows are affected by the configuration of sources, sinks, and models of third-party libraries.
@@ 9. Conclusion
We presented WhyFlow, a tool for end-user debugging for taint analysis. Through speculative analysis, WhyFlow allows developers to ask “why”, “why-not”, and “what-if” questions about the analysis configuration, and is able to visualize multiple, interconnected flows on a graphical view. WhyFlow enables sensemaking of taint analysis and helps users identify root causes of unexpected flows or missing flows. Our user study confirms that WhyFlow helps users provide 21% more correct answers to questions about taint analysis, while experiencing a lower cognitive load. These results suggest that WhyFlow’s inquiry-based approach empowers developers to analyze the outputs of taint analysis.
Data Availability
The replication package and study data are available at https://github.com/UCLA-SEAL/WhyFlow.
Acknowledgements.
This work is supported by the National Science Foundation under grant numbers 2426162, 2106838, and 2106404. It is also supported in part by funding from Amazon and Samsung. We thank the anonymous reviewers for their constructive feedback that helped improve the work.